# LLM Test Generation via Iterative Hybrid Program Analysis

### v1

@@ 4 Class<?> type = null;
  • A fully automated test generation tool, Panta, that supports both generating new test classes and augmenting existing tests.
@@ 6       $CYC_{m}$= computeCyclomaticComplexity($CFG_{m}$)
7       $paths$ = ExplorePaths($CFG_{m}$)
8       methodDict[method] = ($CYC_{m}$, $paths$)
@@ 3.1. Static Path Approximation
Each $CFG_{m}$ is a fully directed graph where code blocks correspond to nodes, and control flow determines the directed edges. We compute the cyclomatic complexity (CYC) (McCabe, 1976) for each method (Line 1), defined as #$\#\text{edges}-\#\text{nodes}+2$. For instance, in Figure 3, we illustrate the $CFG_{m}$ for the parsePattern method from Figure 1. We use the line numbers from Figure 1 to represent nodes, where lines 2 to 5 and lines 9 and 10 form basic code blocks that are grouped into single nodes. Circles represent nodes without branches, while branch nodes are marked as pentagons. The directed edges indicate the control flow between nodes. To ensure a structured representation, we introduce two additional nodes: a start node (1:start) representing the method entry point and a dummy end node (21:end) to unify all execution paths into a single exit. These start and end nodes are depicted as octagons. Excluding these additional nodes, the parsePattern method consists of 12 nodes and 15 edges, with a cyclomatic complexity (CYC) value of 5. We then extract paths using the ExplorePaths in Algorithm 1 (Lines 11), which is based on breadth-first search (BFS). The algorithm collects the shortest set of paths, ensuring each includes at least one unique edge. Exploration terminates once all edges have been visited. By employing ExplorePaths, we guarantee that for each method under test, the number of extracted paths equals its CYC.
@@ 4.1. Experimental Setup
Subject Programs. We evaluate Panta using real-world subjects from Defects4J (Just et al., 2014). Specifically, we selected the latest fixed version of each subject program from Defects4J v2.0.1. To maintain consistency, we exclude three outdated projects—Chart, Mockito, and Closure—as they do not use Maven and rely on deprecated dependencies. The remaining 14 subjects are compatible with Java 8, Maven (v3.6.3), and JUnit (v4.13.2), and they integrate with JaCoCo (v0.8.11) (geeksforgeeks, 2025a) for code coverage measurement and Pitest (v1.17.0) (Coles et al., 2016) for mutation score evaluation.
To evaluate coverage for complex code, rather than generating tests for all classes, we selected non-abstract, public classes that contain at least one method with a cyclomatic complexity greater than 10 (McCabe, 1976; Sanusi et al., 2020). Table 1 presents the selected subjects. The column Project denotes the class name along with its version number from Defects4J. The column #Class represents the number of classes for which we generate tests, while #MUTs counts the total number of methods under test within the selected classes, for each project. In total, we evaluate 130 classes containing 2,971 methods under test (MUTs). Additionally, we report the maximum and median cyclomatic complexity values of the methods under test in the columns $\text{CYC}_{max}$, $\text{CYC}_{med}$, respectively.
@@
LLM Selection. We select the latest open-source model, Meta Llama 3.3 70B (Llama, 2025), as our primary model for evaluation. For RQ3, to assess the impact of different models, we evaluate Panta on three additional models: one open-source model, Mistral Large 2 (Mistral, 2025) from Mistral AI, and two proprietary cost-efficient models—GPT-4o Mini from OpenAI (OpenAI, 2025) and Claude 3.5 Haiku from Anthropic (Anthropic, 2025).
Panta supports both augmenting existing test classes and generating new ones. To ensure a fair comparison, we evaluate Panta in the scenario where tests are generated for a new class without any existing test cases across all research questions.
@@ 4.2. RQ1: Generated Test Quality Comparison
The latest test generation techniques relevant to Panta include HITS (Wang et al., 2024a) and SymPrompt (Ryan et al., 2024). HITS decomposes focal methods into slices and generates tests slice by slice, while SymPrompt extracts path constraints as a list of branch conditions for each method under test and generates tests path by path. Neither approach leverages dynamic code coverage data and both operate statically without guidance on where additional tests are needed. Consequently, HITS and SymPrompt are limited to generating new test classes rather than augmenting existing ones. To differentiate Panta with them, Panta adopts an iterative workflow, guiding the LLM to generate tests for selected paths based on both dynamic and static program analysis information. Our approach works for both scenarios to augment existing tests and add new test classes.
@@ 5. Discussion
Limitations. Since our approach focuses on analyzing one class at a time, its effectiveness diminishes for classes that rely on inheritance. These classes often override parent class methods, resulting in the absence of explicit method implementations, which reduces the effectiveness of control flow analysis. In the future, we plan to extend our work to incorporate cross-class context to address this limitation. Regarding test repair, a common limitation across different LLMs is their inability to effectively fix runtime errors, such as AssertionError caused by failed test assertions. In contrast, LLMs perform better in resolving compilation errors. This indicates that LLMs lack the capability to fully understand project-specific execution logic, as they struggle to infer runtime behaviors and expected outcomes. Regarding test effectiveness, we observe that the average mutation score across different LLMs consistently falls below 50%, indicating that less than half of the introduced mutants are detected. This suggests that LLM-generated tests still have significant room for improvement in enhancing fault detection capability.
Application. The tests generated by Panta operate under the assumption that they are created within a regression testing setting, where the correctness of the underlying focal method implementation is presumed. As Panta supports both augmenting existing test classes and generating new ones, it can be integrated with any test generation tool to further improve code coverage in their generated test suites.
@@ 6. Threats To Validity
Internal Validity. The effectiveness of Panta depends on the capabilities of the LLM used, and results may vary if more advanced models are introduced in the future. To mitigate this, we evaluate Panta using four recently released LLMs, including two open-source and two proprietary models. Another potential issue is the relationship between code coverage and test effectiveness, as high coverage does not necessarily imply meaningful assertions. To address this, we include mutation score in our evaluation to assess fault-detection capability alongside traditional coverage metrics. Additionally, Panta integrates Comex (Das et al., 2023) to generate control flow graphs for path extraction. Any bugs or limitations in Comex could potentially influence our results. To mitigate this risk, we verify the extracted paths during the implementation phase to ensure their correctness before they are used for test generation.
@@
Reproducibility. As this work is still under review, we will make Panta’s implementation and dataset publicly available soon.

### v2

@@ Abstract.
journalyear: 2026copyright: rightsretainedconference: 2026 IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilbooktitle: 2026 IEEE/ACM 48th International Conference on Software Engineering (ICSE ’26), April 12–18, 2026, Rio de Janeiro, Brazildoi: 10.1145/3744916.3764553isbn: 979-8-4007-2025-3/26/04ccs: Software and its engineering Software testing and debugging
@@ 4 Class<?> type = null;
  • A fully automated test generation tool, Panta, that supports both generating new test classes and augmenting existing tests. Panta is publicly available at (Panta, 2025).
@@ 6  $CYC_{m}$= computeCyclomaticComplexity($CFG_{m}$)
7  pathspaths = ExplorePaths($CFG_{m}$)
8  methodDict[method] = ($CYC_{m}$, pathspaths)
@@ 3.1. Static Path Approximation
Each $CFG_{m}$ is a fully directed graph where code blocks correspond to nodes, and control flow determines the directed edges. We compute the cyclomatic complexity (CYC) (McCabe, 1976) for each method (Line 1), defined as $\text{edges}-\#\text{nodes}+2$. For instance, in Figure 3, we illustrate the $CFG_{m}$ for the parsePattern method from Figure 1. We use the line numbers from Figure 1 to represent nodes, where lines 2 to 5 and lines 9 and 10 form basic code blocks that are grouped into single nodes. Circles represent nodes without branches, while branch nodes are marked as pentagons. The directed edges indicate the control flow between nodes. To ensure a structured representation, we introduce two additional nodes: a start node (1:start) representing the method entry point and a dummy end node (21:end) to unify all execution paths into a single exit. These start and end nodes are depicted as octagons. Excluding these additional nodes, the parsePattern method consists of 12 nodes and 15 edges, with a cyclomatic complexity (CYC) value of 5. We then extract paths using the ExplorePaths in Algorithm 1 (Lines 11), which is based on breadth-first search (BFS). The algorithm collects the shortest set of paths, ensuring each includes at least one unique edge. Exploration terminates once all edges have been visited. By employing ExplorePaths, we guarantee that for each method under test, the number of extracted paths equals its CYC.
@@ 4.1. Experimental Setup
Subject Programs. We evaluate Panta using real-world subjects from Defects4J (Just et al., 2014). Specifically, we selected the latest fixed version of each subject program from Defects4J v2.0.1. To maintain consistency, we exclude three outdated projects—Chart, Mockito, and Closure—as they do not use Maven and rely on deprecated dependencies. The remaining 14 subjects are compatible with Java 8 or higher, Maven (v3.6.3), and JUnit (v4.13.2). They are further integrated with JaCoCo (v0.8.11) (geeksforgeeks, 2025a) for code coverage measurement and Pitest (v1.17.0) (Coles et al., 2016) for mutation score evaluation.
To evaluate coverage on complex code, we did not generate tests for all classes; instead, we selected non-abstract, public classes that contain at least one method with cyclomatic complexity greater than 10 (McCabe, 1976; Sanusi et al., 2020). Our process began by automatically selecting all public, non-abstract classes from the 14 Defects4J projects that include at least one non-private method exceeding this threshold. We then manually excluded two types of outliers: (1) classes with extreme complexity (e.g., cyclomatic complexity ¿ 40) caused by large switch-statement patterns, and (2) classes with minimal explicit control flow due to heavy reliance on inheritance. These exclusions ensured that our evaluation remained focused on complex yet analyzable control-flow structures. Table 1 presents the selected subjects. The column Project denotes the class name along with its version number from Defects4J. The column #Class represents the number of classes for which we generate tests, while #MUTs counts the total number of methods under test within the selected classes, for each project. In total, we evaluate 130 classes containing 2,971 methods under test (MUTs). Additionally, we report the maximum and median cyclomatic complexity values of the methods under test in the columns $\text{CYC}_{max}$, $\text{CYC}_{med}$, respectively. More details about individual classes are publicly available in our repository (Panta, 2025).
@@
LLM Selection. We select the latest open-source model, Meta Llama 3.3 70B (Llama, 2025), as our primary model for evaluation. For RQ3, to assess the impact of different models, we evaluate Panta on three additional models: one open-source model, Mistral Large 2 (Mistral, 2025) from Mistral AI, and two proprietary cost-efficient models—GPT-4o Mini from OpenAI (OpenAI, 2025) and Claude 3.5 Haiku from Anthropic (Anthropic, 2025). All LLMs evaluated support a context window of 128k tokens. The model parameters are held constant across all four LLMs used in our evaluation, with max_tokens set to 4096 for output generation and temperature set to 0.2.
Panta supports both augmenting existing test classes and generating new ones. To ensure a fair comparison, we evaluate Panta in the scenario where tests are generated for a new class without any existing test cases across all research questions. Given the low temperature, the generation process is largely deterministic, with only minor fluctuations. As a result, we perform a single run per class and report the average results across all classes for each project.
@@ 4.2. RQ1: Generated Test Quality Comparison
Search-based tools such as EvoSuite (Fraser and Arcuri, 2011) have already been extensively studied in comparison with LLM-based test generation (Wang et al., 2024a; Chen et al., 2024; Ryan et al., 2024; Rao et al., 2023; Deljouyi et al., 2024). Prior results demonstrate that LLM-based tools have advanced beyond search-based approaches by achieving higher average line and branch coverage, although EvoSuite remains competitive on certain projects. Moreover, search-based generated tests are often difficult to interpret and lack meaningful assertions compared to those produced by LLM-based methods (Rao et al., 2023; Deljouyi et al., 2024). Because our contribution lies in advancing LLM-based test generation, we therefore focus our comparisons on other LLM-based tools. The latest test generation techniques relevant to Panta include HITS (Wang et al., 2024a) and SymPrompt (Ryan et al., 2024). HITS decomposes focal methods into slices and generates tests slice by slice, while SymPrompt extracts path constraints as a list of branch conditions for each method under test and generates tests path by path. Neither approach leverages dynamic code coverage data and both operate statically without guidance on where additional tests are needed. Consequently, HITS and SymPrompt are limited to generating new test classes rather than augmenting existing ones. To differentiate Panta with them, Panta adopts an iterative workflow, guiding the LLM to generate tests for selected paths based on both dynamic and static program analysis information. Our approach works for both scenarios to augment existing tests and add new test classes.
@@ 5. Discussion
Limitations. Regarding the sufficiency of source code context, Panta analyzes one class at a time, capturing constructors and internal methods but omitting external dependencies across classes. This design is less effective for classes that rely heavily on inheritance, where key behaviors are defined through method overrides in parent classes. In such cases, the absence of explicit method implementations reduces the effectiveness of control flow analysis. However, since our study examines whether hybrid program analysis improves coverage for complex methods and all evaluated prompting strategies operate in a class-level context, the RQ2 ablation study - conducted under identical conditions - demonstrates a clear improvement with hybrid analysis, validating its effectiveness even within this limited scope. In the future, we plan to extend our approach to incorporate cross-class context to further address this limitation. Regarding test repair, a common limitation across different LLMs is their inability to effectively fix runtime errors, such as AssertionError caused by failed test assertions. In contrast, LLMs perform better in resolving compilation errors. This indicates that LLMs lack the capability to fully understand project-specific execution logic, as they struggle to infer runtime behaviors and expected outcomes. Regarding test effectiveness, we observe that the average mutation score across different LLMs consistently falls below 50%, indicating that less than half of the introduced mutants are detected. This suggests that LLM-generated tests still have significant room for improvement in enhancing fault detection capability.
Application. The tests generated by Panta operate under the assumption that they are created within a regression testing setting, where the correctness of the underlying focal method implementation is presumed. As Panta supports both augmenting existing test classes and generating new ones, it can be integrated with any test generation tool to further improve code coverage in their generated test suites. Our evaluation focuses on Java SE projects from the Defects4J benchmark, which are compatible with Java 8 and newer versions. We select this benchmark because it includes projects with rich control-flow complexity, aligning with our goal of improving coverage for hard-to-cover branches. Consequently, Panta is most applicable to Java SE projects with similar structural complexity. Extending the approach to Java EE applications, which pose additional challenges such as dependency injection and container management, remains an important direction for future work.
Runtime. In our evaluation of 2,971 methods across 130 classes, the average runtime for Panta to generate a passing test suite is 2.3 minutes per method, or approximately 53 minutes per class. Runtime varies with class complexity, method count, and project characteristics. For example, in Cli-40f, classes contain an average of 15 methods with a maximum cyclomatic complexity of 11, resulting in 0.8 minutes per method (12 minutes per class). In contrast, in Collections-28f, classes average 44 methods with cyclomatic complexity up to 32, leading to 9.4 minutes per method (111 minutes per class). Large utility classes with extensive iteration and map usage typically lead to longer generation times. The longer runtimes reflect the challenging nature of the selected classes, which include hard-to-cover branches and high structural complexity. Improving runtime efficiency remains an important avenue for future work.
@@ 6. Threats To Validity
Internal Validity. The effectiveness of Panta depends on the capabilities of the LLM used, and results may vary if more advanced models are introduced in the future. To mitigate this, we evaluate Panta using four recently released LLMs, including two open-source and two proprietary models. Another potential issue is the relationship between code coverage and test effectiveness, as high coverage does not necessarily imply meaningful assertions. To address this, we include mutation score in our evaluation to assess fault-detection capability alongside traditional coverage metrics. Additionally, Panta integrates Comex (Das et al., 2023) to generate control flow graphs for path extraction. Any bugs or limitations in Comex could potentially influence our results. To mitigate this risk, we verify the extracted paths during the implementation phase to ensure their correctness before they are used for test generation. Another concern is data contamination, i.e., the possibility that LLMs may have encountered parts of the Defects4J codebase during pretraining. We recognize this as a broader challenge in LLM-based software engineering research. To mitigate its effect and ensure fair comparisons, all evaluated prompting strategies use the same set of projects and the same underlying LLM. Our evaluation therefore reports relative improvements between prompting strategies rather than absolute performance.
@@
Reproducibility. To ensure reproducibility, we have made Panta’s implementation and dataset publicly available (Panta, 2025). Additionally, we provide detailed instructions for replicating our experimental results.